# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ II-B The Characteristics of QC
where if there is no ambiguity in the context, it is said that the measurement is performed in basis |$\ket{m}$; thereby, the projective operator $P_{m}$ can be written as $P_{m}=\ket{m}\bra{m}$ |.
From the preceding exposition, it is evident that generally, post-measurement qubits |$\ket{\phi}$ alter their state to a particular basis |$\ket{m}$ with probability $p(m)$, unless the basis of measurement coincides with |$\ket{\phi}$ and |ϕketsuperscriptitalic-$\ket{\phi^{\bot}}$. This characteristic is in stark contrast to classical bits, which, irrespective of the form of measurement (observation) or the frequency with which measurements are performed, remain in their initial state. For a state such as |$\ket{\phi}=\sum_{i=0}^{2^{n}-1}c_{i}\ket{i}$, executing a projective measurement $P_{i}=\ket{i}\bra{i}$ | will result in |$\ket{i}$ with a probability of $|c_{i}|^{2}$.
• Quantum states cannot be cloned
A notable attribute of quantum states is that only partial information can be gleaned through a single measurement. It is impossible to ascertain the complete information of a quantum state by repetitively cloning and measuring it. A simplified proof supporting this conclusion is provided herein, with additional specifics available in [22].
Assume a unitary operation $U$ is able to replicate quantum state |$\ket{\phi}$ and substitute quantum resource |$\ket{s}$ with |$\ket{\phi}$:
@@
Apply the inner product on both sides of the above two equations, we obtain:
@@
The resulting equation suggests $x=x^{2}$, whereby $x=0$ or $x=1$ ($\bra{\phi}\ket{\psi}=x$). This necessitates that either |$\ket{\phi}=\ket{\psi}$ or |$\ket{\phi}$ and |$\ket{\psi}$ are orthogonal. As a consequence, a quantum cloning device can only clone mutually orthogonal states, thereby proving that a universal quantum cloning machine cannot exist. In essence, a unitary operation that flawlessly duplicates an unknown quantum state does not exist unless the qubit has degenerated into a classical bit.
@@
The most compelling characteristic of QC is its parallel computing capacity, which arises from the superposition and entanglement of quantum states. As per Postulate 2, QC fundamentally revolves around the unitary evolution of quantum states. Mathematically, unitary evolution can be depicted by a unitary matrix $U_{f}$, which adheres to eq. (18),
@@
where $U_{f}^{\dagger}$ signifies the conjugate transpose of $U_{f}$ and $I$ the identity matrix.
Consider a classical computational function $f:x\rightarrow f(x)$. This function $f$ can be transformed into an associated QC function $U_{f}$ in terms of polynomial cost [23]. Thus, the newly converted $U_{f}$ will implement $U_{f}:\ket{x}\rightarrow\ket{f(x)}$. Suppose an $n$-qubit register stores qubits with initial values |$\ket{00\cdots 0}$. Firstly, $H^{\otimes n}$ is applied in eq. (19) to the qubits, consequently generating a superposition state comprising all binary numbers between 00 and $12^{n}-$ 1 with an equal probability of $\frac{1}{2^{n}}$.
@@
Subsequently, $f(x)$ is computed by $U_{f}$ as per eq. 20,
$U_{f}\!\!\left(\!\!\frac{1}{\sqrt{2^{n}}}\!\sum_{x=0}^{2^{n}-1}\ket{x}\!\!\right)\!\!=\!\!\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}U_{f}\ket{x}\!\!=\!\!\frac{1}{\sqrt{2^{n}}}\!\sum_{x=0}^{2^{n}\!-\!1}\ket{f(x)}$ (20)
@@
This feature of QC, often termed quantum parallelism, is the fundamental reason for its superiority over classical computation. Despite quantum parallel operations typically generating a simultaneous output of $2^{n}$ results, it remains a challenge to read all |$\ket{f(x)}$ states in one go. The sole approach to extracting information from quantum states is through measurement, but each measurement only retrieves a fraction of the resultant information and triggers the collapse of the quantum state. To amass all the information pertaining to the final result, repeated measurements at varying bases (typically exponential) are required, thereby necessitating considerable resources.
@@ II-C The Description of Quantum Algorithm
The preceding exposition of quantum parallelism results in distinctly opposing observations. On one hand, quantum parallelism holds the potential to tremendously boost computing efficiency. Conversely, it can be costly to acquire comprehensive and accurate solutions. Thus, it prompts the question, do quantum algorithms truly offer an advantage?
To answer this question, we should investigate quantum algorithms from the perspective of computational theory through more fundamental problems. How can it be ascertained whether a problem can be addressed using a quantum algorithm? How should a quantum algorithm be described? What’s the mechanism for gauging the computational efficiency and the resource utilization of a quantum algorithm?
The exploration of these questions originated with Benioff’s work [24], wherein he initially coined the quantum Turing machine model. This model was later elaborated by Deutsch [25, 26] and Yao [27], who introduced the quantum circuit model, proving its equivalent standing with the Turing model. Subsequent studies by Bernstein [28] and Knill [29] formed the foundation of quantum computational complexity theory, developing links with classical computational complexity. These classic documents provide the answer:
@@
The Church-Turing proposition claims that any deterministic computational process, regardless of its complexity, can be decided by a Turing machine. From an abacus to a supercomputer, any instrument capable of computation can be simulated by a Turing machine. Hence, all computations can be equated to effective simulations on a Turing machine. To determine whether a mathematical problem is solvable depends on two conditions: whether the input of the problem is countable and whether the Turing machine would stop. Besides the well-known ‘halting problem over Turing machines’, several problems remain unsolvable [30], such as ascertaining if two topological spaces are topologically equivalent (homeomorphic). Can these ostensibly solvable, yet inherently unsolvable problems be resolved on a quantum computer? The answer is negative. A quantum Turing machine is incapable of resolving an issue that a Turing machine cannot, implying that from the standpoint of computability, both the quantum Turing machine and the Turing machine have the same ability. Consequently, problems that are unsolvable through classical methods also remain irresolvable by quantum algorithms.
@@
The ability to solve problems and to solve problems efficiently, are distinct notions. Problems that can be efficiently solved refer to issues resolvable through an algorithm that incurs a polynomial cost relative to the problem scale. It is generally agreed that quantum algorithms can efficiently solve some problems that classical computation cannot address efficiently. While this assertion has not been distinctly proven, substantial evidence supports its validity. Shor’s algorithm [31] is a prime example, efficiently solving large integer factorization and discrete logarithm in $\mathcal{O}(poly(n))$, while no efficient classical algorithms exist to date. Quantum Turing machines deal with vectors, whereas Turing machines handle character strings. Additionally, the transition relations are expected to be more plentiful given the larger number of legal quantum gates compared to logic gates. Accordingly, it is widely accepted that the relation $P\subseteq BQP$ is valid, where BQP represents the class of problems that quantum algorithms can solve efficiently.
@@
The Quantum Turing machine is generally cumbersome for describing quantum algorithms and is more suitable for analyzing and positing the inherent relationship between quantum and classical algorithms. Similar to steering clear of high-level applications in assembly language, a quantum circuit stands as a practical tool for the analysis and description of a quantum algorithm. Yao’s work attests to the equivalence of quantum circuits and quantum Turing machines [27], while Barenco et al. furnished a basic method of analyzing quantum circuits and unveiling the elementary gates within them [32]. These elementary gates form universally complete subsets for describing quantum algorithms. In simple terms, a corresponding quantum circuit can always be constructed to analyze and measure the efficiency and costs of even the most complex quantum algorithms by evaluating circuit depth and gate counts.
Besides quantum circuits, two higher-level ways exist for the description of quantum algorithms. The first approach involves the sequence of unitary operators alone, while the latter builds upon the former, utilizing an Oracle (black box) to replace key steps. Relying on an oracle occurs due to two potential reasons: firstly, the oracle might be trivial, and secondly, theory suggests that an efficient routine should exist to perform these steps, which are currently not a concern. In terms of the number of Oracle accesses, there is another way, quantum query complexity [33, 34], to analyze the complexity of a quantum algorithm.
Even though varied ways of describing quantum algorithms are commonplace in literature, this survey doesn’t strictly discriminate between them due to their inherent equivalency to one another. Still and all, we need to note that generating corresponding quantum circuits (gates) from unitary operators is not trivial. First, the popular superconducting quantum hardware has a sparse near-neighbor connectivity topology and needs to use swap gates to exchange non-adjacent qubits. How to layout the circuit, thereby reducing swap gates, is crucial to the performance. Second, the native gates supported by varied hardware are tremendously different; which gates are adopted to construct quantum circuits will also affect the performance. In fact, this issue, known as “quantum circuit compilation” is challenging and has been proven to be NP-hard [35].
Given that quantum circuits and quantum gates are fundamental for describing and running quantum algorithms, we will proceed to discuss the most frequently used elementary quantum gates in the subsequent section.
@@
Quantum algorithms can be viewed as unitary evolutions of a predetermined initial state within a quantum system. Keeping in mind that unitary operations are synonymous with quantum circuits, decomposing a quantum circuit into elementary gates essentially equates to segmenting complex unitary evolution into fundamental operations. A unified evaluation criterion, specifically a uniform set of elementary gates, streamlines the process of comparing and evaluating the efficiency and costs associated with various quantum algorithms. Conventionally, the elementary gates of quantum circuits are accounted for by single and double qubit gates. Importantly, the set of elementary gates is not static; a significant amount of variation can occur depending on the specific hardware support. Subsequently, we explore some commonly recognized elementary gates.
@@
$\begin{split}X&|0\rangle=|1\rangle\\ X&|1\rangle=|0\rangle\\ X&|\phi\rangle=X(\alpha|0\rangle+\beta|1\rangle)=\alpha|1\rangle+\beta|0\rangle\end{split}$ (21)
@@
$\begin{split}Z&|0\rangle=|0\rangle\\ Z&|1\rangle=-|1\rangle\\ Z&|\phi\rangle=Z(\alpha|0\rangle+\beta|1\rangle)=\alpha|0\rangle-\beta|1\rangle\end{split}$ (22)
@@
$Y=iXZ=i(|1\rangle\langle 0|-|0\rangle\langle 1|)$, $Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$
@@
$\begin{split}Y&|0\rangle=i|1\rangle\\ Y&|1\rangle=-i|0\rangle\\ Y&|\phi\rangle=Y(\alpha|0\rangle+\beta|1\rangle)=i(\alpha|1\rangle-\beta|0\rangle)\end{split}$ (23)
X, Y, and Z are three important single qubit gates, also often represented as Pauli matrices: $\sigma^{x}$, $\sigma^{y}$, and $\sigma^{z}$. These three matrices are not only unitary but also Hermitian, and their eigenvectors form the coordinate axes on the Bloch sphere of a single qubit. They are often used to construct X-, Y-, and Z-axis rotation gates, which are the building blocks for parametric quantum circuits.
@@
$\begin{split}R_{x}(\theta)&=e^{-i\theta X/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}X\\ R_{y}(\theta)&=e^{-i\theta Y/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}Y\\ R_{z}(\theta)&=e^{-i\theta Z/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}Z\end{split}$ (24)
@@
$\begin{split}H|0\rangle&=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle)=|+\rangle\\ H|1\rangle&=\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle)=|-\rangle\\ H|\phi\rangle&\!=\!H(\alpha|0\rangle\!+\!\beta|1\rangle)\!\!=\!\!\frac{1}{\sqrt{2}}(\alpha\!+\!\beta)|0\rangle\!+\!\frac{1}{\sqrt{2}}(\alpha\!-\!\beta)|1\rangle\end{split}$ (25)
$H$ gate has a broad application in generating quantum circuits. For example, acting on |$\ket{00\cdots 0}$ to generate the initial state for most quantum algorithms.
We need to note that the bit flip gate (X) is drawn differently in various literature (Fig. 2a and Fig. 2b), but they both represent the same thing.
@@
Double qubit gates represent unitary operations encompassing two qubits. Prominent among them are the controlled unitary gates, where one qubit is the control end, and the other serves as the controlled qubit. The value of the control qubit, consequently, prescribes whether the unitary operation acts upon the controlled qubit.
In this scenario, we focus on the widely employed CNOT gate, which is alternatively known as the quantum Xor gate. This gate is associated with two inputs: |$\ket{x}$ and |$\ket{y}$, with |$\ket{x}$ in the control position and |$\ket{y}$ as the controlled entity. The semantics attached to the CNOT gate are as follows:
@@
$CNOT=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes X$ (26)
@@
$\begin{split}CNOT|00\rangle=|00\rangle\qquad CNOT|01\rangle=|01\rangle\\ CNOT|10\rangle=|11\rangle\qquad CNOT|11\rangle=|10\rangle\end{split}$ (27)
@@
$\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$ (28)
@@
Figure 3: CNOT gate
@@
It has been established that any unitary matrix within a $d$-dimensional Hilbert space can be decomposed into a product of two-level unitary matrices. Furthermore, it has been elucidated that a combination of single-qubit gates and a CNOT gate can facilitate any two-level unitary operation [36]. Consequently, without a doubt, any treble qubit gate can be simplified by employing a single-qubit gate and a dual-qubit gate. However, a unique treble qubit gate, known as the Toffoli gate, commonly features as a basic module within quantum circuits and usually doesn’t undergo further simplification:
The Toffoli gate comes with three inputs: |$\ket{x}$, |$\ket{y}$, and |$\ket{z}$, where the control qubits are |$\ket{x}$ and |$\ket{y}$, and |$\ket{z}$ operates as the controlled qubit. The semantics of the Toffoli gate are provided below:
@@
Therefore, $Toffli:|x,y,z\rangle\rightarrow|x,y,(x\land y)\oplus z\rangle$, the defination is:
$Toffli=|0\rangle\langle 0|\otimes I\otimes I+|1\rangle\langle 1|\otimes CNOT$ (29)
@@
$\begin{split}Toffli|000\rangle=|000\rangle\qquad Toffli|001\rangle=|001\rangle\\ Toffli|010\rangle=|010\rangle\qquad Toffli|011\rangle=|011\rangle\\ Toffli|100\rangle=|100\rangle\qquad Toffli|101\rangle=|101\rangle\\ Toffli|110\rangle=|111\rangle\qquad Toffli|111\rangle=|110\rangle\end{split}$ (30)
@@
$\begin{pmatrix}1&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1\\ 0&0&0&0&0&0&1&0\end{pmatrix}$ (31)
Toffoli gate representation is shown in fig. 4a. In particular, when $|z\rangle=|0\rangle$, the Toffoli gate can be exploited as a quantum ‘AND’ gate (fig. 4b).

### v3

@@ II-B The Characteristics of QC
If there is no ambiguity in the context, it is said that the measurement is performed in basis |$\ket{m}$; thereby, the projective operator $P_{m}$ can be written as $P_{m}=\ket{m}\bra{m}$ |.
From the preceding discussion, it is clear that, in general, post-measurement qubits |$\ket{\phi}$ collapse to a particular basis state |$\ket{m}$ with probability $p(m)$, unless the measurement basis aligns with |$\ket{\phi}$ and its orthogonal complement |ϕketsuperscriptitalic-$\ket{\phi^{\bot}}$. This behavior contrasts with classical bits, which retain their initial state regardless of the form or frequency of observations (measurements). For a quantum state such as |$\ket{\phi}=\sum_{i=0}^{2^{n}-1}c_{i}\ket{i}$, performing a projective measurement $P_{i}=\ket{i}\bra{i}$ | will yield the outcome |$\ket{i}$ with a probability of $|c_{i}|^{2}$.
• Quantum states can not be cloned.
A key property of quantum states is that only limited information can be extracted from a single measurement. It is fundamentally impossible to fully determine a quantum state by repeatedly cloning and measuring it, due to the no-cloning theorem. A simplified proof of this principle is presented here, with further details available in [20].
Assume a unitary operation $U$ is able to replicate |$\ket{\phi}$ and substitute resource |$\ket{s}$ with |$\ket{\phi}$:
@@
Apply the inner product on both sides of eqs. (15) and (16), we obtain:
@@
The resulting equation suggests $x=x^{2}$, where $x=0$ or $x=1$ ($\bra{\phi}\ket{\psi}=x$). This implies that either |$\ket{\phi}=\ket{\psi}$ or |$\ket{\phi}$ and |$\ket{\psi}$ is orthogonal. Consequently, a quantum cloning device can only clone mutually orthogonal states, thereby proving that a universal quantum cloning machine cannot exist. In essence, a unitary operation that perfectly replicates an unknown quantum state does not exist unless the qubit has effectively collapsed into a classical bit.
@@
The defining feature of QC is its ability to perform parallel computation, enabled by the principles of superposition and entanglement of quantum states. According to Postulate 2, QC operates on the unitary evolution of quantum states. Mathematically, this evolution is represented by a unitary matrix $U_{f}$, which satisfies the condition in eq. (18),
@@
where $U_{f}^{\dagger}$ signifies the conjugate transpose of $U_{f}$ and identity matrix $I$.
Consider a classical computational function $f:x\rightarrow f(x)$. This function $f$ can be transformed into a corresponding QC function $U_{f}$ with polynomial computational cost [21]. The transformed function $U_{f}$ will then operate as $U_{f}:\ket{x}\rightarrow\ket{f(x)}$. Assume an $n$-qubit register holds qubits with initial values |$\ket{00\cdots 0}$. Initially, the Hadamard transform $H^{\otimes n}$ is applied, as shown in eq. (19), resulting in a superposition state of all binary numbers between 00 and $12^{n}-$ 1, each occurring with equal probability of $\frac{1}{2^{n}}$.
@@
Subsequently, $f(x)$ is computed by $U_{f}$ in eq. 20,
$\!U_{f}\!\!\left(\!\!\frac{1}{\sqrt{2^{n}}}\!\sum_{x=0}^{2^{n}-1}\ket{x}\!\!\right)\!\!=\!\!\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}U_{f}\ket{x}\!\!=\!\!\frac{1}{\sqrt{2^{n}}}\!\sum_{x=0}^{2^{n}\!-\!1}\ket{f(x)}.$ (20)
Eq. (20) demonstrates that $U_{f}$ performs all computations of |$\ket{f(x)}$ for inputs ranging from |$\ket{00\cdots 0}$ to |$\ket{11\cdots 1}$. In contrast, on a classical computer, calculating $f(x)$ for every $x$ from $00\cdots 0$ to $11\cdots 1$ requires either $2^{n}$ sequential loops or $2^{n}$ parallel processors.
This capability of QC, referred as quantum parallelism, strengthens its advantage over classical computation. Although quantum parallel operations can generate $2^{n}$ results simultaneously, reading all |$\ket{f(x)}$ states in one step presents a significant challenge. Information from quantum states can only be extracted via measurement, but each measurement reveals only partial information and collapses the quantum state. To gather the full spectrum of information on the final outcome, multiple measurements in different bases (often exponentially large in number) are needed, which requires substantial computational resources.
3) Entanglement
Entanglement can be viewed as a specialized form of superposition, exhibiting properties of quantum parallelism, since quantum parallelism originates from the superposition of quantum states. For example, consider the Bell state $\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})$, which represents a superposition of the states |$\ket{00}$ and |$\ket{11}$. When inputs are |$\ket{00}$ and |$\ket{11}$, a suitable unitary operation can process these states simultaneously, demonstrating quantum parallelism.
@@
Similarly, the state $\frac{1}{\sqrt{3}}(\ket{00}+\ket{10}+\ket{11})$ is also entangled since it can not be expressed as a separable tensor product. However, its entanglement is weaker than that of the Bell state $\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})$. This difference in entanglement can be quantitatively measured using entanglement entropy, as defined in eq. (21),
$S(\rho)=-tr(\rho^{A}\log\rho^{A})=-\sum_{i}\lambda_{i}\log\lambda_{i},$ (21)
where $\rho^{A}$ is reduced density operator of subsystem A of a composite system $\rho^{AB}$. $\rho^{A}$ is defined as a partial trace over system B in eq. (22):
[…]
where |$\ket{a_{1}}$ and |$\ket{a_{2}}$ are vectors of subsystem A, and |$\ket{b_{1}}$ and |$\ket{b_{2}}$ are vectors of subsystem B.
Simple calculations reveal that the entanglement entropy of the Bell state is 1, while the state $\frac{1}{\sqrt{3}}(\ket{00}+\ket{10}+\ket{11})$ has an entanglement entropy of approximately 0.17, significantly lower than that of the Bell state.
Entanglement plays a crucial role in quantum communication protocols, such as super-dense coding and quantum teleportation. However, in QML, the effect of entangled data on prediction error presents a dual nature, contingent upon the number of allowed measurements. When a sufficient number of measurements are available, increasing the entanglement of the training data consistently reduces prediction error or lowers the required size of the training set to achieve a comparable error rate. On the other hand, when the number of measurements is limited, highly entangled data may result in increased prediction errors [22].
@@ II-C The Description of Quantum Algorithm
The preceding discussion on quantum parallelism presents two contrasting views. On one hand, quantum parallelism has the potential to significantly enhance computational efficiency. On the other hand, obtaining complete and precise solutions can be resource-intensive. This leads to a fundamental question: do quantum algorithms truly provide an advantage?
To address the problem, we need to examine quantum algorithms from the perspective of computational theory and basic problem structures. How can one determine if a problem is suitable for a quantum algorithm? How a quantum algorithm is described? And what methods can be used to evaluate its computational efficiency and resource usage?
This exploration began with Benioff’s foundational work [23], where he introduced the quantum Turing machine model. This model was later expanded by Deutsch [24] and Yao [25], who introduced the quantum circuit model, showing that it is equivalent to the Turing machine model. Further work by Bernstein [26] and Knill [27] established the theoretical foundation of quantum computational complexity, linking it to classical complexity theory. These classic studies provide a key insight:
@@
According to the Church-Turing thesis, any deterministic computational process, no matter how complex, can be represented by a Turing machine. This means that all computations can be effectively simulated by a Turing machine, whether performed by an abacus or a supercomputer. The solvability of a problem, therefore, depends on whether the input is countable and whether the Turing machine halts. Aside from the well-known halting problem, several problems remain unsolvable [28], such as determining whether two topological spaces are homeomorphic.
Are these theoretically solvable but practically are unsolvable problems solved on a quantum computer? The answer is no. A quantum Turing machine can not solve problems that a classical Turing machine can not solve, meaning that in terms of computability, both models are equivalent. Therefore, problems that are unsolvable by classical means remain unsolvable by quantum algorithms as well.
@@
The distinction between solving problems and solving them efficiently is critical. A problem that can be efficiently solved implies that there exists an algorithm capable of addressing the problem with a cost that grows polynomially with the size of the input. It is generally acknowledged that quantum algorithms can efficiently solve certain problems that classical computation can not handle efficiently. Although this claim has not been definitively proven, there is substantial evidence supporting its validity.
Shor’s algorithm [29] serves as a notable example, as it efficiently solves problems like large integer factorization and discrete logarithms in $O(\text{poly}(n))$, while no classical algorithms can achieve this efficiently. The key difference lies in the nature of the computational models. Quantum Turing machines manipulate quantum states (vectors), whereas classical Turing machines operate on character strings. Moreover, quantum systems have a greater variety of allowed operations due to the larger set of valid quantum gates compared to classical logic gates.
Consequently, it is widely accepted that $P\subseteq BQP$, where $BQP$ represents the class of problems that can be efficiently solved by quantum algorithms. This inclusion indicates that quantum computation can address at least some problems more efficiently than classical methods.
@@
The quantum Turing machine, while useful for understanding the theoretical relationship between quantum and classical algorithms, is generally different for practical algorithm description. Like high-level applications abstracted away from assembly language, quantum circuits offer a more practical framework for analyzing and describing quantum algorithms. Yao [25] demonstrates the equivalence between quantum circuits and quantum Turing machines [25]. [30] provides foundational methods for analyzing quantum circuits and identifying their elementary gates. These elementary gates form a universal set for constructing and analyzing quantum algorithms. Consequently, any quantum algorithm, regardless of complexity, can be represented and evaluated in terms of circuit depth and gate count.
In addition to quantum circuits, there are two higher-level methods for describing quantum algorithms. The first method involves a sequence of unitary operators, while the second method builds upon the first by incorporating an oracle (a black box that performs specific functions). The use of an oracle is often justified because it might represent a trivial operation or because an efficient procedure should exist for those operations theoretically, which is not the primary concern here. Quantum query complexity [31] is another approach to analyzing the complexity of a quantum algorithm based on the number of oracle queries.
While various methods for describing quantum algorithms are discussed in the literature, this survey does not strictly differentiate among them due to their inherent equivalency. However, constructing quantum circuits from unitary operators is not straightforward. For example, popular superconducting quantum hardware features sparse near-neighbor connectivity, requiring the use of swap gates to exchange non-adjacent qubits. The efficient layout of the circuit to minimize the use of swap gates is crucial for performance. Additionally, the native gates supported by different hardware platforms vary significantly, impacting performance based on which gates are used to build quantum circuits. This problem, known as “quantum circuit compilation,” has been proven NP-hard [32].
Given that quantum circuits and gates are essential for both describing and executing quantum algorithms, the next section will focus on discussing the most widely used elementary quantum gates.
@@
Quantum algorithms can be conceptualized as unitary evolutions starting from a predetermined initial state within a quantum system. Since unitary operations are essentially quantum circuits, decomposing a quantum circuit into elementary gates is equivalent to breaking down complex unitary evolutions into fundamental operations. A standardized evaluation criterion, using a uniform set of elementary gates, facilitates the comparison and assessment of the efficiency and costs associated with different quantum algorithms. Typically, the elementary gates of quantum circuits include single and two-qubit gates. It is important to note that the set of elementary gates is not fixed. It varies significantly depending on the hardware support. The following section examines some of the widely recognized elementary gates.
@@
$\begin{split}X&|0\rangle=|1\rangle\\ X&|1\rangle=|0\rangle\\ X&|\phi\rangle=X(\alpha|0\rangle+\beta|1\rangle)=\alpha|1\rangle+\beta|0\rangle\end{split}.$ (23)
@@
$\begin{split}Z&|0\rangle=|0\rangle\\ Z&|1\rangle=-|1\rangle\\ Z&|\phi\rangle=Z(\alpha|0\rangle+\beta|1\rangle)=\alpha|0\rangle-\beta|1\rangle\end{split}.$ (24)
@@
$Y=iXZ=i(|1\rangle\langle 0|-|0\rangle\langle 1|)$, $Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$.
@@
$\begin{split}Y&|0\rangle=i|1\rangle\\ Y&|1\rangle=-i|0\rangle\\ Y&|\phi\rangle=Y(\alpha|0\rangle+\beta|1\rangle)=i(\alpha|1\rangle-\beta|0\rangle)\end{split}.$ (25)
$X$, $Y$, and $Z$ gates are fundamental single-qubit gates, frequently represented by the Pauli matrices: $\sigma^{x}$, $\sigma^{y}$, and $\sigma^{z}$. These matrices are not only unitary but also Hermitian, with their eigenvectors defining the coordinate axes on the Bloch sphere of a single qubit. These gates are essential for constructing $X$-, $Y$-, and $Z$-axis rotation gates, which serves as the building blocks for parameterized quantum circuits.
@@
$\begin{split}R_{x}(\theta)&=e^{-i\theta X/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}X\\ R_{y}(\theta)&=e^{-i\theta Y/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}Y\\ R_{z}(\theta)&=e^{-i\theta Z/2}=cos\frac{\theta}{2}I-isin\frac{\theta}{2}Z\end{split}.$ (26)
@@
$\begin{split}H|0\rangle&=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle)=|+\rangle\\ H|1\rangle&=\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle)=|-\rangle\\ H|\phi\rangle&\!=\!H(\alpha|0\rangle\!+\!\beta|1\rangle)\!\!=\!\!\frac{1}{\sqrt{2}}(\alpha\!+\!\beta)|0\rangle\!+\!\frac{1}{\sqrt{2}}(\alpha\!-\!\beta)|1\rangle\end{split}.$ (27)
Hadamard gate ($H$) is widely used in generating quantum circuits, such as creating the initial state |$\ket{00\cdots 0}$ for many quantum algorithms. It is important to note that the bit-flip gate ($X$) in Figs. (2a) and (2b) can be depicted differently in various sources. Actually, these representations are equivalent.
@@
Double qubit gates perform unitary operations involving two qubits. Among these, the controlled unitary gates are notable, where one qubit acts as the control qubit and the other as the target (controlled) qubit. The state of the control qubit determines whether the unitary operation is applied to the target qubit.
Here, we focus on the widely used CNOT gate in Fig. (3), also known as the quantum XOR gate. This gate operates on two qubits: |$\ket{x}$ as the control qubit and |$\ket{y}$ as the target qubit. The CNOT gate operates as follows:
@@
$CNOT=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes X.$ (28)
@@
$\begin{split}CNOT|00\rangle=|00\rangle\qquad CNOT|01\rangle=|01\rangle\\ CNOT|10\rangle=|11\rangle\qquad CNOT|11\rangle=|10\rangle\end{split}.$ (29)
@@
$\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}.$ (30)
@@
Figure 3: CNOT gate.
@@
It has been established that any unitary matrix within a $d$-dimensional Hilbert space can be decomposed into a product of two-level unitary matrices. Furthermore, it has been shown that a combination of single-qubit gates and a CNOT gate can facilitate any two-level unitary operation [33]. Consequently, any three-qubit gate can be simplified using single-qubit gates and two-qubit gates. However, a specific three-qubit gate, known as the Toffoli gate, widely features as a fundamental module within quantum circuits and usually does not undergo further simplification:
The Toffoli gate has three inputs: |$\ket{x}$, |$\ket{y}$, and |$\ket{z}$, where |$\ket{x}$ and |$\ket{y}$ are control qubits and |$\ket{z}$ is the target qubit. The semantics of the Toffoli gate are as follows:
@@
Therefore, $\text{Toffli}:|x,y,z\rangle\rightarrow|x,y,(x\land y)\oplus z\rangle$, the definition is:
$\text{Toffli}=|0\rangle\langle 0|\otimes I\otimes I+|1\rangle\langle 1|\otimes CNOT.$ (31)
@@
$\begin{split}\text{Toffli}|000\rangle=|000\rangle\qquad\text{Toffli}|001\rangle=|001\rangle\\ \text{Toffli}|010\rangle=|010\rangle\qquad\text{Toffli}|011\rangle=|011\rangle\\ \text{Toffli}|100\rangle=|100\rangle\qquad\text{Toffli}|101\rangle=|101\rangle\\ \text{Toffli}|110\rangle=|111\rangle\qquad\text{Toffli}|111\rangle=|110\rangle\end{split}.$ (32)
@@
$\begin{pmatrix}1&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1\\ 0&0&0&0&0&0&1&0\end{pmatrix}.$ (33)
Toffoli gate representation is shown in Fig. (4a). In particular, when $|z\rangle=|0\rangle$, the Toffoli gate can be exploited as a quantum ‘AND’ gate in Fig. (4b).